OpenStudy (anonymous):

g(x)=4/(8-3x) find the domain of the function

OpenStudy (anonymous):

all real numbers except for 8/3 because that would make the denominator 0 and the fraction would be unsolvable.

OpenStudy (anonymous):

you have to figure out when the denominator equals 0

OpenStudy (anonymous):

\[g(x) = \frac{4}{8 - 3x}\] The denominator cannot be zero because it is undefined so we have to check what value will make the denominator zero so \[8 - 3x = 0\] Solve for x and you get \[x = \frac{8}{3}\]. X cannot be 8/3 because it will make the denominator zero which is undefined. So the domain is \[(-\infty,\frac{8}{3})\cup(\frac{8}{3},\infty)\]